### v2

# FIAN/TD/04–25

Abstract
A new interpretation of Dirac singletons [1], i.e.,  free conformal fields in $d$ dimensions, as relativistic fields in a $d+1$-dimensional space-time with cosmological constant, that differs from the Flato-Fronsdal dipole construction in $AdS_{d+1}$ [2], is proposed. The $d+1$-dimensional field is described at the level of both equations and Lagrangian. It forms an infinite-dimensional representation of the $d+1$-dimensional Lorentz group that relates fields at different space-time points. The associated well-known fact is that singleton cannot be localized at a point in ${d+1}$ dimensions, hence being unobservable via local scattering/radiation phenomena in the Standard Model ($d=3$). On the other hand, that singleton respects ${d+1}$ dimensional relativistic symmetries makes it possible to introduce its interactions with gravity and other relativistic fields in $d+1$ dimensions. It is speculated that the presence of singleton in a four-dimensional field theory with non-zero cosmological constant (dark energy) can be relevant to the dark matter phenomenon and baryon asymmetry generation.
1 Introduction
Aleksei Starobinsky was one of the world leaders in the field of quantum gravity and cosmology (see, e.g., [3]). He passed away tragically in a few days because of the Covid 19 decease. I had a privilege to know Aleksei personally though our scientific interests were different enough. Alexei was a deep knowledgable person whose interests spread far beyond the scientific research. In this paper, dedicated to the memory of Aleksei, I wish to make a comment that may indicate some convergence between our seemingly different fields.
The aim of this letter is to point out that there exists an object (field) very different from the fields usually used in the Standard Model (SM), that may, in principle, affect general picture of the origin of some of phenomena in high-energy physics, cosmology and astrophysics. The main actor of this letter is the so-called singleton discovered by Dirac in 1963 in [1] [^1] The name was given even earlier in the math literature [4] to emphasize its simplicity from the representation theory perspective. as a specific branch of the solutions of a certain wave equation, that survives at infinity of $AdS_{4}$. There are two singleton fields of boson and fermion types often called $Rac$ and $Di$, respectively [5] (see [6, 7] for reviews). Later on it was realized that $Rac$ and $Di$ and their supermultiplets considered in the context of supergravity $S^{7}$ compactifications, supermembrane and higher-spin theory, for instance, in [8]-[15] are nothing else as, respectively, free massless scalar and spinor fields in the 2+1 dimensional space presently identified with the boundary of $AdS_{4}$ within the paradigm of holographic correspondence [16]-[18], which interpretation is particularly relevant in the context of higher-spin holography [19, 20].
In this letter we focus on the two issues.
Firstly, we interpret singletons as relativistic fields in $(A)dS_{d+1}$ providing for them Lorentz covariant field equations and actions which, to the best of our knowledge, were not available before. This is achieved within the unfolded dynamics approach (see [21] for an elementary review and more references). The proposed formulation differs from the Flato-Fronsdal dipole one [2] given in terms of a certain fourth-order field equation in $AdS_{d+1}$. In particular, it is free from the huge gauge symmetry of [2] eliminating local degrees of freedom in $AdS_{d+1}$. On the other hand, kinematically our formulation has some similarities with the interpretation of holography proposed in the framework of the nonlinear realisation approach in [22, 23] (though not at the Lagrangian level).
Secondly, we speculate on the possible implications of the presence of the singleton field in the context of SM and Gravity. Let us stress that our proposal differs from seemingly analogous ideas of Flato and Fronsdal engineering usual fields, starting from the electromagnetic one [24]-[27] (and references therein), as composites of singletons. Instead we treat singletons democratically with all other fields of SM and Gravity as independent fields.
In particular, we speculate that singleton, that exists in presence of dark energy, can be relevant to the dark matter problem as well as some other physical phenomena including the baryon asymmetry. It is important that though the genuine singleton is defined in $AdS_{4}$ our construction is applicable to $dS_{4}$ as well. Since singleton makes sense in presence of a nonzero cosmological constant (=dark energy), this idea somehow unifies dark matter with dark energy.
To put it short, our goal is to draw attention to the fact that, apart from usual local relativistic fields, there may exist exotic relativistic objects that are not local in the usual sense. Their characteristic feature is that they form an infinite-dimensional representation of the Lorentz group. A related phenomenon that singleton cannot be localised at a point in the $3d$ space of $4d$ space-time is known from the very first singleton papers [2, 7]. As a result, from the $4d$ space-time perspective it is nowhere (= everywhere). This means in particular that singleton does not affect the local scattering and radiation SM processes. On the other hand, that singletons form representations of the Lorentz algebra makes it possible to introduce their gravitational interaction within Cartan formalism. As argued in the paper, singleton can contribute to collective physical phenomena beyond the standard collider physics.
The paper is organized as follows. In Section 2 we sketch main ideas of the unfolded dynamics approach underlying our construction. The mechanism allowing to formulate the same dynamics in spaces of different dimensions is recollected in Section 3. Unfolded formulation of the conformal scalar and spinor in $d$ dimensions is presented in Sections 4 and 5, respectively, at the both off-shell and on-shell levels, including conformal invariant Lagrangians. Extension of the singletons to $(A)dS_{d+1}$ is presented in Section 6. Spinor formulation of the on-shell $3d$ singletons and their $4d$ extensions is presented in Section 7. In Section 8 we speculate on the potential role of singletons for the resolution of some problems in SM and related aspects of cosmology and astrophysics. Main results and further research directions are summarised in Section 9.
2 Unfolded dynamics
The unfolded form of dynamical equations is a generalization of the first-order form of ordinary differential equations
$\dot{q}^{i}(t)=\varphi^{i}(q(t))\,$ (2.1)
resulting from the replacement of the time derivative by the de Rham derivation,
$\frac{\partial}{\partial t}\to{\rm d}=\theta^{n}\partial_{n}\,,\qquad\theta^{n}\theta^{m}=-\theta^{m}\theta^{n}\,,\qquad dx^{n}\equiv\theta^{n}\,,$
and the dynamical variables $q^{i}$ by a set of differential forms $W^{\Omega}(\theta,x)$
$q^{i}(t)\rightarrow W^{\Omega}(\theta,x)=\sum_{p=0}\theta^{n_{1}}\ldots\theta^{n_{p}}W^{\Omega}_{n_{1}\ldots n_{p}}(x)\,,$
that allows one to reformulate a system of partial differential equations in the first-order form
${\rm d}W^{\Omega}(\theta,x)=G^{\Omega}(W(\theta,x))\,.\qquad$ (2.2)
Here $G^{\Omega}(W)$ are some functions of the “supercoordinates” $W^{\Omega}$,
$G^{\Omega}(W)=\sum_{n}f^{\Omega}{}_{\Lambda_{1}\ldots\Lambda_{n}}W^{\Lambda_{1}}\ldots W^{\Lambda_{n}}\,.$
Since ${\rm d}^{2}=0$, at $d>1$ the functions $G^{\Lambda}(W)$ have to obey the compatibility conditions
$G^{\Lambda}(W)\frac{\partial G^{\Omega}(W)}{\partial W^{\Lambda}}\equiv 0\,.$ (2.3)
Let us stress that these are conditions on the functions $G^{\Lambda}(W)$ rather than on $W^{\Omega}$.
The idea of the unfolded formulation was put forward in [28] where it was realized that the full system of nonlinear equations for massless higher-spin gauge fields can be searched in the form (2.2) as a deformation of the free unfolded equations.
As a consequence of the compatibility conditions (2.3) the system (2.2) is invariant under the gauge transformation
$\delta W^{\Omega}={\rm d}\varepsilon^{\Omega}+\varepsilon^{\Lambda}\frac{\partial G^{\Omega}(W)}{\partial W^{\Lambda}}\,,$ (2.4)
where the gauge parameter $\varepsilon^{\Omega}(x)$ is a $(p_{\Omega}-1)$-form if $W^{\Omega}$ was a $p_{\Omega}$-form. Strictly speaking, this is true for the class of universal unfolded systems in which the compatibility conditions (2.3) hold independently of the dimension $d$ of space-time, i.e.,  (2.3) should be true disregarding the fact that any $(d+1)$-form is zero. This is the case in all unfolded systems considered in the higher-spin literature including this paper.
As shown in [29], the variety of invariant functionals associated with the unfolded equations (2.2) is described by the cohomology of the operator
$Q=G^{\Omega}\frac{\partial}{\partial W^{\Omega}}\,,$ (2.5)
that obeys
$Q^{2}=0$ (2.6)
as a consequence of (2.3). As a result, the unfolded equations can be written in the Hamiltonian-like form
${\rm d}F(W)=Q(F(W))\,,\qquad\forall F(W)\,.$ (2.7)
By virtue of (2.7), $Q$-closed $p$-forms $L_{p}(W)$ are ${\rm d}$-closed, giving rise to the gauge invariant functionals
$S=\int_{\Sigma^{p}}L_{p}\,.$ (2.8)
(See [29] for more detail and examples.)
A particular example of an unfolded system is provided by the zero-curvature (Maurer-Cartan) equations. Namely, let $h$ be a Lie algebra with the basis $\{T_{\alpha}\}$. Let $\omega=\omega^{\alpha}T_{\alpha}$ be a one-form valued in $h$. For $G(\,\omega)=-\omega\omega:=-\frac{1}{2}\omega^{\alpha}\omega^{\beta}[T_{\alpha},T_{\beta}]$ equation (2.2) with $W=\omega$ reads as [^2] The exterior product wedge symbol is omitted in this paper since all products are automatically exterior due to the presence of anticommuting $\theta$.
${\rm d}\omega+\omega\omega=0\,.$ (2.9)
Relations (2.3) and (2.4) amount, respectively, to the usual Jacobi identity for $h$ and gauge transformation for $\omega$.
If the set $W^{\alpha}$ contains some $p$-forms ${\cal C}^{i}$ (e.g. zero-forms) and the functions $G^{i}$ are linear in $\omega$ and ${\cal C}$,
$G^{i}=-\omega^{\alpha}(T_{\alpha})^{i}{}_{j}{\cal C}^{j}\,,$ (2.10)
then (2.3) implies that the matrices $(T_{\alpha})^{i}{}_{j}$ form some representation $T$ of $h$, acting in a space $V$ where ${\cal C}^{i}$ is valued. The associated equation (2.2) is a covariant constancy condition
$D_{\omega}{\cal C}=0$ (2.11)
with $D_{\omega}\equiv{\rm d}+\omega$ being a covariant derivative in the $h$-module $V$.
The zero-curvature equations (2.9) usually describe background geometry in a coordinate independent fashion. For instance, let $h$ be the Poincaré algebra with the gauge fields
$\omega(x)=e^{n}(x)P_{n}+\frac{1}{2}\omega^{nm}(x)M_{nm}\,,$ (2.12)
where $P_{n}$ and $L_{nm}$ are generators of translations and Lorentz transformations with $e^{n}(x)$ and $\omega^{nm}(x)$ identified with the frame one-form and Lorentz connection, respectively (fiber Lorentz vector indices $m,n\ldots$ run from 0 to $d-1$ and are raised and lowered by the flat Minkowski metric). It is well known that the zero-curvature condition (2.9) for the Poincaré algebra describes Minkowski geometry in a coordinate-independent way.
By choosing a different Lie algebra $h$ one can describe a different background like, e.g., anti-de Sitter for $h=o(d-1,2)$ or conformally flat for $h=o(d,2)$. The covariant constancy equation (2.11) then describes $h$-invariant linear equations in a chosen background.
The unfolded formulation has a number of remarkable properties starting from its general applicability: every system of partial differential equations can be reformulated in the unfolded form. Due to the exterior algebra formalism, the system is coordinate independent.
Local degrees of freedom are represented by the subset of zero-forms $C^{I}(x_{0})\in\{W^{\Omega}(x_{0})\}$ at any $x=x_{0}$. This is analogous to the fact that $q^{i}(t_{0})$ describe degrees of freedom in the first-order form of ordinary differential equations. In field-theoretic models, the zero-forms $C^{I}(x_{0})$ realize an infinite-dimensional module dual to the space of single-particle states of the system in question [30]. This space is analogous to the phase space in the Hamiltonian dynamics.
3 Space-time metamorphoses and holography
Unfolded dynamics exhibits independence of the “world-volume” space-time with coordinates $x$. Instead, geometry is encoded by the functions $G^{\Omega}(W)$ in the “target space” with fields $W^{\Omega}$ as local coordinates. Since the universal unfolded equations make sense in any space-time independently of a particular realization of the de Rham derivation ${\rm d}$, one is free to extend space-time by additional coordinates $z^{u}$,
${\rm d}W^{\Omega}(x)=G^{\Omega}(W(x))\,,\quad x\rightarrow X=(x,z)\,,\quad{\rm d}_{x}\rightarrow{\rm d}_{X}={\rm d}_{x}+{\rm d}_{z}\,,\quad{\rm d}_{z}=dz^{u}\frac{\partial}{\partial z^{u}}\,.$ (3.13)
Locally, unfolded equations reconstruct the $X$-dependence in terms of values of the fields $W^{\Omega}(X_{0})=W^{\Omega}(x_{0},z_{0})$ at any $X_{0}$. Clearly, to take $W^{\Omega}(x_{0},z_{0})$ in space $M_{X}$ with coordinates $X_{0}$ is the same as to take $W^{\Omega}(x_{0})$ in the space $M_{x}\subset M_{X}$ with coordinates $x$.
Generally, unfolding can be interpreted as some sort of covariant twistor transform [21, 31]
Here $W(Y|x)$ are functions on the “correspondence space” $C$ with local coordinates $Y,x$. The space-time $M$ has local coordinates $x$. The twistor space $\mathbf{T}$ has local coordinates $Y$ the expansion over which generates the components $W^{\Omega}(x)$ with various $\Omega$ (for examples see Sections 4, 5, 7).
Unfolded equations reconstruct the dependence of $W(Y|x)$ on $x$ in terms of the function $W(Y|x_{0})$ on $\mathbf{T}$ at some $x_{0}$. The restriction of $W(Y|x)$ or some its $Y$-derivatives to $Y=0$ gives dynamical fields $\omega(x)$ in $M$ which, in the on-shell case, solve their dynamical field equations. Hence, similarly to the Penrose transform (see [21] and references therein), unfolded equations map functions on $\mathbf{T}$ to solutions of the dynamical field equations in $M$.
In these terms, the holographic duality can be interpreted as the duality between different space-times $M$ that can be associated with the same twistor space. The problem becomes most interesting provided that there is a nontrivial vacuum connection along the additional coordinates $z$. This is in particular the case of $AdS/CFT$ correspondence where the conformal flat connection at the boundary is extended to the flat $AdS$ connection in the bulk with $z$ being a Poincaré coordinate. This mechanism has a number of interesting applications. In particular, in [30] it was applied to the description of all $4d$ massless fields in terms of a single scalar in the ten-dimensional space identified by Fronsdal with Lagrangian Grassmannian in [32] (see also [33, 34]).
Conventional holography [16]-[18] is based on the isomorphism of the boundary conformal group $O(d,2)$ with the symmetry of $AdS_{d+1}$. Generators of the conformal algebra obey the relations
$[D\,,P_{a}]=-P_{a}\,,\qquad[D\,,K^{b}]=K^{b}\,,\qquad[D\,,L_{ab}]=0\,,$ (3.14)
$[P_{a}\,,K_{b}]=2L_{ab}-2\eta_{ab}D\,,$ (3.15)
$[L_{ab}\,,P_{c}]=\eta_{bc}P_{a}-\eta_{ac}P_{b}\,,\qquad[L_{ab}\,,K_{c}]=\eta_{bc}K_{a}-\eta_{ac}K_{b}\,,$ (3.16)
$[L_{ab}\,,L_{cd}]=\eta_{bc}L_{ad}-\eta_{ac}L_{bd}-\eta_{bd}L_{ac}+\eta_{ad}L_{bc}\,.$ (3.17)
Let $M^{d}$ be a $d$–dimensional conformally flat space-time with local coordinates ${\bf x}$ and some $o(d,2)$ connection $w_{{\bf x}}({\bf x})=w^{A}_{{\bf x}}T_{A}$ obeying the flatness conditions
${\rm d}_{\bf x}w_{{\bf x}}({\bf x})+w_{{\bf x}}({\bf x})w_{{\bf x}}({\bf x})=0\,,\qquad{\rm d}_{\bf x}:=d{\bf x}^{n}\frac{\partial}{\partial{\bf x}^{n}}\,.$ (3.18)
A particular flat connection, that corresponds to Cartesian coordinates in $M^{d}$, is
$w_{{\bf x}}({\bf x})=d{\bf x}^{a}P_{a}\,.$ (3.19)
The dilatation generator $D$ induces standard $\mathbb{Z}$ grading on $o(d,2)$,
$[D\,,T_{A}]=\Delta(T_{A})T_{A}\,$ (3.20)
with $\Delta(T_{A})$ being the conformal dimension of $T_{A}$,
$\Delta(L)=0\,,\qquad\Delta(D)=0\,,\qquad\Delta(K)=1\,,\qquad\Delta(P)=-1\,.$ (3.21)
Let us now introduce an additional coordinate $z$ and differential $dz$ so that $x=({\bf x},z)$ be local coordinates of $AdS_{d+1}$. A conformally foliated connection $W(x)$ of $AdS_{d+1}$ can be introduced as follows. The components of the connection with differentials $d{\bf x}$ are
$W_{{\bf x}}^{A}(x)T_{A}=z^{\Delta(T_{A})}w^{A}_{{\bf x}}({\bf x})T_{A}\,,$ (3.22)
while the only nonzero $dz$ component of the connection is associated with the dilatation generator $D$, having the form
$W_{z}(x)D=-z^{-1}dzD\,.$ (3.23)
Clearly, so defined connection $W(x)$ is flat in (a local chart of) $AdS_{d+1}$. Poincaré coordinates result from this construction applied to the connection (3.19).
Analogously, unfolded equations
${\mathcal{D}}_{\bf x}C_{i}({\bf x})=0\,,\qquad{\mathcal{D}}_{\bf x}:={\rm d}_{\bf x}+W^{A}_{\bf x}T_{A}$ (3.24)
in $M^{d}$ for a set of fields $C_{i}({\bf x})$ carrying conformal weights $\Delta_{i}$ extend to the fields
$C_{i}(x)=z^{\Delta_{i}}C_{i}({\bf x})\,$ (3.25)
and equations
${\mathcal{D}}_{x}C_{i}(x)=0\,,\qquad{\mathcal{D}}_{x}:={\rm d}_{x}+W^{A}_{x}T_{A}\,.$ (3.26)
It is important to note that if a system was off-shell in $M^{d}$ this is not so in the extended $d+1$-dimensional space. Indeed, the dependence on the additional coordinate $z$ is determined by (3.26) in terms of that on ${\bf x}$ as is most obvious from (3.25). This means that the field in $AdS_{d+1}$ obeys some differential equation, that determines its $z$-dependence.
To identify the $d+1$-dimensional space with (a local chart of) $AdS_{d+1}$ it suffices to redefine $o(d,2)$ generators as
${\mathcal{P}}_{\nu}=\big{(}(P_{a}+\lambda^{2}K_{a})\,,2\lambda D\big{)}\,,\qquad{\mathcal{M}}_{\nu\mu}=\big{(}L_{ab},\frac{1}{2\lambda}(P_{a}-\lambda^{2}K_{a})\delta_{\nu}^{d},-\frac{1}{2\lambda}(P_{b}-\lambda^{2}K_{b})\delta_{\mu}^{d}\big{)}$ (3.27)
with $a,b=(0\,,\ldots,d-1)$ $\nu,\mu=(0\,,\ldots,d)$, interpreting ${\mathcal{P}}_{\nu}$ and ${\mathcal{M}}_{\nu\mu}$ as $AdS_{d+1}$ translation (transvection) and Lorentz generators, respectively. The respective connections are
$W=h^{\nu}{\mathcal{P}}_{\nu}+\frac{1}{2}\omega^{\nu\mu}{\mathcal{M}}_{\nu\mu}\,.$ (3.28)
In particular, this means that
$E^{a}=\frac{1}{2}(h^{a}+2\lambda\omega^{ad})\,,$ (3.29)
where $E^{a}$ is the $d$-dimensional vielbein rescaled in accordance with (3.22) (the index $d$ in (3.29) indicates the direction along $z$). The dimensionful parameter $\lambda$ is related to the cosmological constant $\Lambda$,
$\Lambda=-\#\lambda^{2}\,$ (3.30)
with some positive number ##\##. This implies that $\lambda$ is real and pure imaginary in the $AdS_{d+1}$ and $dS_{d+1}$ cases, respectively. Naively, this suggests that the construction is not working in the de Sitter space. In fact, this is not necessarily true and, moreover, as discussed below, specificities of the $dS$ case may be of relevance to the baryon asymmetry problem.
According to the Flato-Fronsdal theorem [5], boundary conformal currents are dual to fields in the bulk $AdS_{d+1}$. In other words, free relativistic fields in the bulk are associated with the bilinear currents on the boundary, the fact underlying the Klebanov-Polyakov HS holographic conjecture [20]. Here we consider an opposite situation: starting from the boundary conformal field we will see what is its bulk dual. Our construction differs from the other holographic treatments of singletons (see e.g. [6, 35, 36]) based on the dipole singleton description of [2]. The output is interesting both formally and, hopefully, physicswise shedding more light on what are bulk duals of the free conformal boundary fields, i.e.,  singletons.
4 Conformal scalar in any $d$ within unfolded formalism
Singleton $Rac$ is a massless conformal scalar field in any dimension $d$. In the unfolded dynamics approach it is described as follows [37]. Let $C(y|{\bf x})$ be a zero-form, that depends on the space-time coordinates ${\bf x}^{n}$ and auxiliary variables $y^{n}$ ($n=0,\ldots d-1$). Consider unfolded equations of the form
${\rm d}_{\bf x}C(y|{\bf x})+d{\bf x}^{n}\frac{\partial}{\partial y^{n}}C(y|{\bf x})=0\,.$ (4.31)
Clearly, this equation relates the coefficients $C_{a_{1}\ldots a_{n}}({\bf x})$ of the expansion
$C(y|{\bf x})=\sum_{n=0}^{\infty}\frac{1}{n!}C_{a_{1}\ldots a_{n}}({\bf x})y^{a_{1}}\ldots y^{a_{n}}$ (4.32)
to higher ${\bf x}^{a}$ derivatives of the ground component $C({\bf x})$,
$C_{a_{1}\ldots a_{n}}({\bf x})=(-1)^{n}\partial_{a_{1}}\ldots\partial_{a_{n}}C({\bf x})\,,\qquad\partial_{a}:=\frac{\partial}{\partial{\bf x}^{a}}\,,$ (4.33)
$C({\bf x}):=C(0|{\bf x})\,$ (4.34)
identified with the scalar field. The system (4.31) is off-shell, imposing no differential conditions on $C({\bf x})$. To put the system on shell of a massless field it suffices to constrain $C(y|{\bf x})$ by the condition
$\Box_{y}C(y|{\bf x})=0\,,\qquad\Box_{y}:=\eta^{ab}\frac{\partial^{2}}{\partial y^{a}\partial y^{b}}\,.$ (4.35)
The system (4.31) is equivalent to
${\mathcal{D}}_{\bf x}C(y|{\bf x})=0\,,$ (4.36)
where
${\mathcal{D}}_{\bf x}:={\rm d}_{\bf x}+e^{a}P_{a}+f_{a}K^{a}+\frac{1}{2}\omega^{ab}L_{ab}+bD$ (4.37)
is the covariant derivative of the conformal algebra $o(d,2)$ with the generators $P_{a}$ for translations, $K^{a}$ for special conformal transformations, $L_{ab}$ for Lorentz transformations and $D$ for dilatations. The particular flat connection used in (4.31) is
$e^{a}=d{\bf x}^{a}\,,\qquad\omega^{ab}=0\,,\qquad b=0\,,\qquad f_{a}=0\,.$ (4.38)
In terms of $y^{a}$, the conformal generators, that obey (3.14)-(3.17), are realized as
$P_{a}=\frac{\partial}{\partial y^{a}}\,,\qquad L_{ab}=y_{a}\frac{\partial}{\partial y^{b}}-y_{b}\frac{\partial}{\partial y^{a}}\,,\qquad D=y^{a}\frac{\partial}{\partial y^{a}}+\Delta\,,$ (4.39)
$K_{a}=y^{2}\frac{\partial}{\partial y^{a}}-2y_{a}y^{b}\frac{\partial}{\partial y^{b}}-2\Delta y_{a}\,,$ (4.40)
where $\Delta$ is a number (conformal weight). The system (4.36) along with the consistency condition
${\mathcal{D}}_{\bf x}^{2}=0$ (4.41)
equivalent to the flatness condition (3.18) forms an unfolded system invariant under the gauge conformal transformations (2.4). Any choice of a particular flat connection $e^{a}$, $f_{a}$, $\omega_{ab}$, $b$ in (4.37) restricts the local conformal transformations (2.4) to the global ones. (Analogously the choice of Minkowski metric restricts diffeomorphisms to global Poincaré transformations. For more detail see [21]). This proves global conformal invariance of the system (4.36).
Note that the construction of this section is extendable to $(A)dS_{d}$ as well as to any other conformally flat background by the appropriate choice of the flat connection of the conformal group. In particular, conformal field theories in $AdS_{d}$ and their holographic aspects were discussed in [30, 38, 39, 36].
Next one observes that if $C(Y|{\bf x})$ obeys the constraint (4.35), it is obeyed by $T_{A}C(Y|{\bf x})$ for all conformal algebra generators $T_{A}$ provided that
$\Delta=\frac{d}{2}-1\,,$ (4.42)
which is the canonical dimension of a massless scalar in $d$ dimensions. This implies that the constraint (4.35) is preserved by the equation (4.36). Since the system (4.35), (4.36) is equivalent to the massless Klein-Gordon equation, this in turn proves conformal invariance of the latter.
From the representation theory perspective this implies that the module $V^{tr}$ generated by the conformal generators from a constant in $y^{a}$ only contains $y$–traceless polynomials $f(y)$ obeying $\Box_{y}f(y)=0\,.$ To see this it is instructive to check that $K^{a}K_{a}$ applied to a constant yields zero for $K^{a}$ (4.40) and $\Delta$ (4.42). Hence, $V^{tr}$ is a submodule of the module $V$ of all polynomials, $V^{tr}\subset V$, while traceful polynomials in $y^{a}$ are in the factor module $V/V^{tr}$.
Now we observe that the coefficients $C_{a_{1}\ldots a_{n}}({\bf x})$ in the expansion (4.32) are in the dual module $V^{*}$ with respect to the action of the covariant derivative ${\mathcal{D}}$ (4.36), (4.37). This means that $(V/V^{tr})^{*}$ is a submodule of $V^{*}$ while $(V^{tr})^{*}$ is a factor-module. As a result, there are two components of $C(y)$ the covariant derivative of which does not contain the gauge field $f^{a}$ of special conformal transformations. One is the vacuum (lowest weight) component $C({\bf x})$, while another is the singular vector (synonymous to be annihilated by $K^{a}$) associated with the trace component
$C^{\prime}({\bf x}):=C^{a}{}_{a}({\bf x})\,.$ (4.43)
This implies that to prove conformal invariance of any functional built from $C({\bf x})$ and $C^{\prime}({\bf x})$ it suffices to check its invariance under the action of the parabolic subalgebra generated by $P_{a}$, $L_{ab}$ and $D$. (For more detail on the derivation of general conformal invariant equations from the representation theory of the conformal group see [40] and references therein.)
The conformal invariant Lagrangian for a scalar field is a $d$-form
$L^{Rac}=\frac{1}{2}\epsilon_{a_{1}\ldots a_{d}}e^{a_{1}}({\bf x})\ldots e^{a_{d}}({\bf x})C({\bf x})C^{\prime}({\bf x})\,,$ (4.44)
where $e^{a}$ is a vielbein one-form in $d$ dimensions. It is easy to see that $L^{Rac}$ is $Q$-closed with respect to $Q$ (2.5). Indeed, the special conformal gauge field does not appear in $QL$ since it is absent in $de^{a}$, ${\mathcal{D}}(C)$ and ${\mathcal{D}}(C^{\prime})$. Lorentz connection $\omega^{ab}$ cancels because $L$ is Lorentz invariant. Analogously, $b$ cancels because the Lagrangian has proper scaling dimension due to (4.42). Finally, the contribution of $e^{b}$ cancels because of antisymmetrization over $d+1$ indices $a=0,1,\ldots d-1$ due to the exterior product (i.e.,  $\theta$–dependence) of the one-forms $e^{a}$.
The fields $C(y|{\bf x})$ still obey unfolded equations (4.36), that are off-shell just expressing higher components $C_{a_{1}\ldots a_{n}}({\bf x})$ via derivatives of $C({\bf x})$ according to (4.33), imposing no differential conditions on the latter. In particular, they imply that $C^{\prime}({\bf x})=\Box_{\bf x}C({\bf x})$. The Lagrangian field equations then imply the massless equation $\Box_{\bf x}C(x)=0$.
5 Conformal spinor in any $d$ within unfolded formalism
Conformal spinor $Di$ is described analogously to scalar $Rac$. Consider unfolded equations of the form
${\rm d}_{\bf x}C_{\alpha}(y|{\bf x})+d{\bf x}^{a}\frac{\partial}{\partial y^{a}}C_{\alpha}(y|{\bf x})=0\,$ (5.45)
with a spinor index $\alpha$. This equation relates the coefficients $C_{\alpha,a_{1}\ldots a_{n}}({\bf x})$ of the expansion
$C_{\alpha}(y|{\bf x})=\sum_{n=0}^{\infty}\frac{1}{n!}C_{\alpha,a_{1}\ldots a_{n}}({\bf x})y^{a_{1}}\ldots y^{a_{n}}$ (5.46)
to higher derivatives in ${\bf x}^{a}$,
$C_{\alpha,a_{1}\ldots a_{n}}({\bf x})=(-1)^{n}\partial_{a_{1}}\ldots\partial_{a_{n}}C_{\alpha}({\bf x})\,,$ (5.47)
where $C_{\alpha}({\bf x})$ is the ground component of $C_{\alpha}(y|{\bf x})$,
$C_{\alpha}({\bf x}):=C_{\alpha}(0|{\bf x})\,$ (5.48)
identified with the genuine spinor field. As such, the system is off-shell, imposing no differential conditions on $C_{\alpha}({\bf x})$.
To put the system on shell of a massless field it suffices to impose the constraint on $C(y|{\bf x})$
$\gamma^{a}{}_{\alpha}{}^{\beta}\frac{\partial}{\partial y^{a}}C_{\beta}(y|{\bf x})=0\,,$ (5.49)
where $\gamma^{a}$ are gamma-matrices,
$[\gamma^{a}\,,\gamma^{b}]=2\eta^{ab}Id\,.$ (5.50)
The system (5.45) is the particular case of
${\mathcal{D}}_{\bf x}C_{\alpha}(y|{\bf x})=0\,$ (5.51)
for ${\mathcal{D}}_{\bf x}$ of the form (4.37), (4.38) with the conformal generators
$P_{a}=\frac{\partial}{\partial y^{a}}\,,\qquad L_{ab}=y_{a}\frac{\partial}{\partial y^{b}}-y_{b}\frac{\partial}{\partial y^{a}}+\frac{1}{4}[\gamma_{a}\,,\gamma_{b}]\,,\qquad D=y^{a}\frac{\partial}{\partial y^{a}}+\Delta\,,$ (5.52)
$K_{a}=y^{2}\frac{\partial}{\partial y^{a}}-2y_{a}y^{b}\frac{\partial}{\partial y^{b}}-2\Delta y_{a}+\frac{1}{2}y^{b}[\gamma_{b}\,,\gamma_{a}]\,.$ (5.53)
It is not hard to see that, for
$\Delta=\frac{d-1}{2}\,,$ (5.54)
which is a canonical dimension of the massless spinor, the action of $K^{a}\gamma_{a}$ on a $y^{a}$-independent element yields zero. This implies that $\gamma^{a}$–transversal polynomials form a submodule $V^{\gamma tr}$ of $V^{sp}$ of all polynomials (5.46). Analogously to the scalar case, this has a consequence that the space of spinor fields (5.46) possesses a lowest weight vector $C_{\alpha}({\bf x})$ identified with the dynamical spinor field and a singular vector $C^{\prime}_{\alpha}({\bf x})$ associated with the Dirac operator,
$C_{\alpha}({\bf x}):=C_{\alpha}(0|{\bf x})\,,\qquad C^{\prime}_{\alpha}({\bf x}):=\gamma^{a}{}_{\alpha}{}^{\beta}C_{\beta a}(0|{\bf x})\,.$ (5.55)
The conformally invariant Lagrangian has the form analogous to (4.44),
$L^{Di}=\frac{1}{2}\epsilon_{a_{1}\ldots a_{d}}e^{a_{1}}({\bf x})\ldots e^{a_{d}}({\bf x})\bar{C}^{\alpha}({\bf x})C^{\prime}_{\alpha}({\bf x})\,,$ (5.56)
where $\bar{C}^{\alpha}({\bf x})$ is the Dirac conjugated spinor. Clearly, in the Cartesian coordinate system (4.38), Eq. (5.56) yields usual Dirac Lagrangian for the spinor field $C_{\alpha}({\bf x})$ in $d$-dimensions.
6 Extension to $(A)dS_{d+1}$
Now we are in a position to extend the $d$-dimensional singleton systems to $AdS_{d+1}$. To this end we replace equations (4.36), (5.51) by analogous equations
$D_{x}C(x)=0\,,\qquad D_{x}C_{\alpha}(x)=0\,,\qquad D_{x}:={\rm d}_{x}+W$ (6.57)
with $W$ (3.28). The conformal generators have the form (4.39), (4.40) and (5.52), (5.53) in the scalar and spinor cases, respectively.
The $AdS_{d+1}$ invariant Lagrangians still have the form (4.44) for scalar and (5.56) for spinor but now being $d$-forms in $(A)dS_{d+1}$ with the fields $C$ and $C^{\prime}$ rescaled by (3.25) and $E^{a}$ (3.29),
$L^{Rac}=\frac{1}{2}\epsilon_{a_{1}\ldots a_{d}}E^{a_{1}}(x)\ldots E^{a_{d}}(x)C(x)C^{\prime}(x)\,,$ (6.58)
$L^{Di}=\frac{1}{2}\epsilon_{a_{1}\ldots a_{d}}E^{a_{1}}(x)\ldots E^{a_{d}}(x)\bar{C}^{\alpha}(x)C^{\prime}_{\alpha}(x)\,.$ (6.59)
These Lagrangians are closed and, as a consequence of the general properties of the unfolded equations, invariant up to exact forms (i.e.,  total derivatives) under the symmetries (2.4) that leave invariant the background connections, that is global $(A)dS$ symmetries. However, now one has to take into account that equations (6.57) are no longer off-shell reconstructing the dependence of one of the coordinates, namely $z$, in terms of the others. This is the reason why the seemingly non-invariant form of (6.58), (6.59) in view of (3.29) still respects Lorentz covariance in $AdS_{d+1}$. In other words, the $d+1$-dimensional Lorentz transformations act on the singleton nonlocally relating fields $\phi({\bf x},z)$ at different $z$. As a result, being a local field in $d$ dimensions, from the $d+1$ perspective it is nowhere (equivalently, everywhere).
As mentioned in Section 3, the parameter $\lambda$, that enters the $(A)dS$ connection by virtue of (3.27), (3.28), is real in the $AdS$ case but pure imaginary in $dS$. Naively, this implies that the Lagrangians (6.58) and (6.59) are not Hermitian in the $dS$ case. However, this problem can be resolved by introducing doublets of mutually conjugated fields $C^{\pm}$ and/or $C^{\pm}_{\alpha}$, associated with $\lambda=\pm i\lambda^{\prime}$ with real $\lambda^{\prime}$. This allows one to consider singletons as fields in the de Sitter space tantamount to dark energy [3]. The modes associated with the evolution along $z$ are either increasing or decreasing that is not too surprising in the expansion regime.
7 Spinor formulation for $3d$ singletons
In 2+1 dimensions it is convenient to describe massless fields in terms of spinor indices $\alpha,\beta=1,2$. In these terms, local coordinates are ${\bf x}^{\alpha\beta}={\bf x}^{\beta\alpha}$ and massless field equations are
$\frac{\partial}{\partial{\bf x}^{\alpha\beta}}\frac{\partial}{\partial{\bf x}_{\alpha\beta}}\phi({\bf x})=0$ (7.60)
and
$\frac{\partial}{\partial{\bf x}^{\alpha\beta}}\phi^{\beta}({\bf x})=0$ (7.61)
for scalar $\phi({\bf x})$ and spinor $\phi^{\beta}({\bf x})$, respectively. The unfolded massless field equations are formulated [41] both for $Rac$ and for $Di$ with the aid of auxiliary spinor variables $y^{+}_{\alpha}$, $y^{-}_{\beta}$ obeying
$[y^{-}_{\alpha}\,,y^{+}_{\beta}]=\varepsilon_{\alpha\beta}$
in terms of the Fock module
$|\phi(y^{+}|{\bf x})\rangle=\phi(y^{+}|{\bf x})|0\rangle\,,\qquad y^{-}_{\alpha}|0\rangle=0\,,\qquad\phi(y^{+}|x)=\sum_{n=0}^{\infty}\frac{1}{n!}\phi_{\alpha_{1}\ldots\alpha_{n}}(x)y^{+\alpha_{1}}\ldots y^{+\alpha_{n}}\,$ (7.62)
in the form
$D_{\bf x}|\phi(y^{+}|{\bf x})\rangle=0$ (7.63)
with the $o(3,2)\sim sp(4|\mathbb{R})$ covariant derivative
$D_{\bf x}:={\rm d}_{\bf x}+e^{\alpha\beta}({\bf x})y^{-\alpha}y^{-\beta}+\omega^{\alpha\beta}({\bf x})y^{+}_{\alpha}y^{-}_{\beta}+b({\bf x})\{y^{+\alpha}\,,y^{-}_{\alpha}\}+f({\bf x})_{\alpha\beta}y^{+\alpha}y^{+\beta}\,$ (7.64)
for some flat connection obeying $D_{\bf x}^{2}=0$. Cartesian coordinate system results from the flat connection
$e^{\alpha\beta}({\bf x})=d{\bf x}^{\alpha\beta}\,,\qquad\omega^{\alpha\beta}=0\,,\qquad b=0\,,\qquad f_{\alpha\beta}=0\,.$ (7.65)
To describe singleton as a $4d$ field it suffices to add an additional coordinate and extend the flat $o(3,2)$ connection to the four-dimensional space. Identifying the additional coordinate with the Poincaré coordinate $z$ we set
${\bf x}^{\alpha\beta}\to x^{\alpha{\dot{\alpha}}}=({\bf x}^{\alpha{\dot{\alpha}}},-\frac{i}{2}\epsilon^{\alpha{\dot{\alpha}}}z^{-1})\,.$
The unfolded field equations in $AdS_{4}$ read as
$D_{x}|\phi\rangle=0\,$ (7.66)
with
$D_{x}={\rm d}_{x}+\frac{i}{z}d{\bf x}^{\alpha\beta}y^{-}_{\alpha}y^{-}_{\beta}-\frac{dz}{2z}y^{-}_{\alpha}y^{+\alpha}\,,\qquad{\rm d}_{x}:=dx^{\alpha{\dot{\beta}}}\frac{\partial}{\partial x^{\alpha{\dot{\beta}}}}$
describing a flat $AdS_{4}$ connection in Poincaré coordinates with
$e^{\alpha{\dot{\alpha}}}=\frac{1}{2z}dx^{\alpha{\dot{\alpha}}}\,,\qquad\omega^{\alpha\beta}=-\frac{i}{4z}d{\bf x}^{\alpha\beta}\,,\qquad\bar{\omega}^{{\dot{\alpha}}{\dot{\beta}}}=\frac{i}{4z}d{\bf x}^{{\dot{\alpha}}{\dot{\beta}}}\,.$
(For more detail see [42].)
Though unfolded equations are formulated in the $4d$ space-time, the dynamical equations are still three-dimensional
$\frac{\partial^{2}}{\partial{\bf x}^{\alpha\beta}\partial{\bf x}_{\alpha\beta}}\phi({\bf x},z)=0$
at every $z$. The equations (7.66) then reconstruct $z$ dependence of $\phi({\bf x},z)$ in terms of $\phi({\bf x},z_{0})$ at any $z_{0}\neq 0$. The system is relativistic in the $4d$ sense since the Lorentz algebra $o(3,1)$ is a subalgebra of $o(3,2)$ as well as of $o(4,1)$. However, the singleton field treated as a relativistic field in the $4d$ space-time is very different from usual local space-time fields since it belongs to an infinite-dimensional $o(3,1)$ module, namely, the Fock module (7.62). This is because the $4d$ Lorentz generators $L^{\alpha\beta}$ and $\bar{L}^{{\dot{\alpha}}{\dot{\beta}}}$ realized as
$L_{\alpha\beta}=\frac{1}{2}(y^{+}_{\alpha}+iy^{-}_{\alpha})(y^{+}_{\beta}+iy^{-}_{\beta})\,,\qquad\bar{L}_{{\dot{\alpha}}{\dot{\beta}}}=\frac{1}{2}(y^{-}_{\dot{\alpha}}+iy^{+}_{\dot{\alpha}})(y^{-}_{\dot{\beta}}+iy^{+}_{\dot{\beta}})$ (7.67)
(see [42]) contain the creation $y^{+}y^{+}$ parts.
To formulate an off-shell Lagrangian description of the $3d$ singleton in terms of spinors one has to introduce along the lines of the $4d$ construction of [43] an additional scalar variable $p$ parameterising the (gamma)traceful components of the off-shell $3d$ fields $(C_{\alpha}({\bf x}))C({\bf x})$. The off-shell extension will be elaborated elsewhere.
8 Singleton as a beyond SM actor
That $4d$ Lorentz algebra acts on the singleton is crucially important making it possible to introduce interaction of the singleton with $4d$ gravity in the Cartan formalism. One can write a Lagrangian as a sum of the singleton Lagrangians with the usual $4d$ matter Lagrangians of the SM, GR, and further extensions,
$L=L^{Rac}+L^{Di}+L^{SM}+L^{GR}+\ldots\,.$ (8.68)
As such, it is a sum of the three-form Lagrangians $L^{Rac}$ and $L^{Di}$ and four-form Lagrangians for the genuine $4d$ local fields contributing to $L^{SM}$ and $L^{GR}$. Let us stress that since the Lagrangians are closed forms, the respective actions are insensitive to the local variations of the three-cycles over which $L^{Rac}$ and $L^{Di}$ are integrated in the action,
$S=\int_{\Sigma_{Rac}^{3}}L^{Rac}+\int_{\Sigma_{Di}^{3}}L^{Di}+\int_{M^{4}}(L^{SM}+L^{GR}+\ldots)\,.$ (8.69)
Because singletons are not local $4d$ fields, their direct scattering effects can unlikely be observable in the collider experiments. This implies that the presence of singletons should not affect the local high-energy $4d$ SM physics, thus avoiding a tension with the available experimental data. The same time, this raises a question whether the singletons may yield observable phenomena whatsoever.
There is an alternative mechanisms that may play a role, however, namely the prominent Flato-Fronsdal theorem [5] stating that the tensor product of two singleton fields (equivalently, their bilocal composits) amounts to the direct sum of $4d$ massless fields including the massless spin two field, i.e.,  graviton, and a massless scalar in the singlet representation with respect to all inner symmetries,
$S\bigotimes S=\sum_{s=0}^{\infty}\phi_{s,m=0}(x)=graviton+singlet\,\,scalar+\ldots\,.$ (8.70)
For the extesion of Flato-Fronsdal theorem to the case with inner symmetries see [44, 45] while its harmonic analysis interpretation was given in [46]. A related idea of bilocal fields was also put forward in [47].
That bilinears of singletons contain graviton and an inert massless scalar may have observable consequences via, e.g., an additionally induced gravitational field as well as invisible (dark) matter. In other words, though the $4d$ singleton fields cannot be localized themselves in some region of a galaxy, their nonlinear combinations can induce localizable fields. Of course, to evaluate such effects one has to introduce and analyse interactions. The first step towards gravitational interaction consists of the covariantization of $D_{x}$ in (7.66) beyond the flat $AdS_{d+1}$ connection. Note, however, that, as usual in the unfolded dynamics, this is only the first step that demands further extension to respect the formal consistency in the sense of (2.3) beyond the linearized approximation. The simplest way towards solution to this problem is probably via construction of the unfolded version of $3d$ conformal gravity theory with $3d$ massless matter. The $4d$ singleton theory then will result via the application of the space-time metamorphoses mechanism of Section 3 to the $3d$ conformal theory. For some progress in this direction within $3d$ conformal higher-spin gravity see [48, 49], while unfolded formulation of conformal geometry was considered in [50]. The construction of the nonlinear $3d$ conformal higher-spin theory interacting with singletons was proposed in the recent papers [51, 52]. The problem is to evaluate the effect of condensation of the gravitational and scalar fields in the galaxy induced by the singleton.
An intriguing group-theoretic singleton phenomenon found in [14] is that the compactification of eleven-dimensional supergravity on the seven sphere leads to singleton representations along with the usual $4d$ relativistic representations in the spectrum. Moreover, in the squashed seven sphere case singletons were argued to be involved into certain higgsing together with the usual $4d$ unitary representations. Since the field-theoretic interpretation of this phenomenon is still not clear it would be interesting to apply the developed unfolded machinery to its further analysis. The formalism developed in this paper may be most appropriate in this respect since unfolded dynamics essentially maps the representation theory to dynamical equations. Also a potentially useful for better understanding of this phenomenon is the coset space approach to $AdS/CFT$ of [23]. It would be interesting to elaborate more on its relation to the unfolded dynamics approach used in this paper.
Note that Flato-Fronsdal theorem suggests that the tensor product of singletons contains massless fields of all spins. This is a group-theoretic fact true at the free field level that respects higher-spin symmetries. In the situation considered in this paper with singletons interacting with the fields of SM and gravity this is no-longer true and the tail of higher-spin fields is anticipated to be deformed to some kind of effective interactions. On the other hand, since lower-spin relativistic and inner symmetries remain unbroken, the graviton and scalar field contributions to (8.70) are anticipated to survive.
It is important to stress that, to be dynamically active, singleton should live in the $(A)dS$ space. In other words the proposed construction can only be working in presence of dark energy. More precisely, it works directly in the $sp(4)\sim o(3,2)$ invariant $AdS_{4}$ space with the negative cosmological constant $\Lambda$ while dark energy associated with the $dS_{4}$ space has opposite sign of $\Lambda$. In $dS_{4}$, the singleton Lagrangians contain imaginary unit via $\lambda:=\sqrt{-\Lambda}$. However, as discussed in Section 6, by an appropriate doubling of fields one can make the model Hermitean. The presence of exponential solutions is natural in the $dS_{4}$ regime of expanding Universe. (Note that singleton in the flat space was argued to be a constant [2, 53, 54].)
The presence of complex coefficients in the singleton equations may be related to such a persisting problem in cosmology and high-energy physics as baryon asymmetry. Indeed, these properties fit the prominent Sakharov necessary conditions [55] for baryon asymmetry (for review see, e.g., [56]). Positive cosmological constant may provide a non-equilibrium regime. Moreover, singletons endowed with appropriate inner structure may induce violation of the baryon number conservation. The presence of complex coefficients in the Lagrangian may induce $CP$ violation. Though, because of its $4d$ nonlocality, the singleton field can hardly be seen in usual scattering phenomena, it can affect the formation of a charged matter via a nonlinear manifestation of the Flato-Fronsdal theorem.
A related comment is that being non-localisable in the $4d$ space such fields behave as background charges in the effective field theory if they have some non-zero VEVs responsible for violation of some global symmetries beyond the list of those manifest in the SM. In that case the proposed idea may have some relation to another Dirac’s idea of evolution of fundamental (cosmological) constants [57] via evolution of the singleton fields.
9 Conclusion
The goal of this paper is to point out that there is an unusual type of $4d$ relativistic field, the Dirac singleton, that exists in presence of cosmological constant (dark energy). The new result of the paper is the formulation of its dynamics directly in the $4d$ space-time in a way free of the necessity of factorisation of the bulk modes as in the dipole approach of [2]. The singleton matter is unusual in the sense that it is $4d$ non-localisable. This is closely related to the fact that singleton forms an infinite-dimensional representation of the $4d$ Lorentz group, that does not mean, however, that singleton has more degrees of freedom than usual relativistic fields associated with finite-dimensional tensor-spinor Lorentz representations. Just other way around, singleton is essentially a $3d$ conformal field of usual type.
In this respect it is interesting to compare the proposed construction with an alternative recent dark matter candidate suggested by Bogomolny [58] based on another nonstandard relativistic matter proposed by Dirac [59]. The difference is that the latter equation describes infinitely many degrees of freedom in four dimension while singleton, being a $3d$ field, carries less than one usual relativistic field. It should be noted that Flato and Fronsdal (see [7] and references therein) suggested to use singleton as a kind of constituent matter with usual relativistic fields realized as its composites. In this paper singleton is described directly in four dimensions as a relativistic field coexisting with other relativistic fields of usual types.
Note that one can treat analogously other conformal fields in appropriate space-time dimensions. For instance, $4d$ spin-one massless field is conformal and hence can be uplifted to a relativistic singleton-type field in $(A)dS_{5}$ as discussed in [60]. More general types of singletons associated with unitary conformal fields in even dimensions can also be considered [61, 62].
Let us stress again that, since the singleton field is nonlocal from the $4d$ perspective, it can hardly be observed via a radiation process but can affect the surrounding gravitational field to be observed via the matter motion, that may induce a dark matter-like effect. Another potential application discussed in Section 8 is for the baryon asymmetry explanation. More generally, singletons may admit some cosmological manifestations affecting the cosmological evolution as well. This is what the scientific interests of Alexei Starobinsky were focused on.
Acknowledgments
I wish to thank Konstantin Alkalaev, Mikhail Alfimov, Andrei Barvinsky, Olga Gelfond, Sergey Godunov, Carlo Iazeolla, Evgeny Ivanov, Andrei Mironov, Nikita Misuna, Roman Nevzorov, Bengt Nilsson, Dmitry Ponomarev, Aleksandr Pukhov, Arkady Tseytlin, Boris Voronov, and Mikhail Vysotsky for useful comments. I am grateful for hospitality to Ofer Aharony, Theoretical High Energy Physics Group of Weizmann Institute of Science where a part of this work has been done. This work was supported by Theoretical Physics and Mathematics Advancement Foundation BASIS Grant No 24-1-1-9-5.